OpenStudy (anonymous):

The hyperbola with center at (5, -6), with vertices at (0, -6) and (10, -6), with (15, 0) a point on the hyperbola.

OpenStudy (anonymous):

Have you determined whether this is an upward-downward opening hyperbola or a left-right opening hyperbola?

OpenStudy (anonymous):

I tried to graph it but the vertices ended up in the same quadrant

OpenStudy (anonymous):

Well, that's fine, because the center is not at the origin. The hyperbola is just translated into the first quadrant. Is it up-down or left-right? Plotting the center will help.

OpenStudy (anonymous):

*fourth quadrant

OpenStudy (anonymous):

left right

OpenStudy (anonymous):

Good. A general equation for a left-right opening parabola centered at the origin with radius \(r\) is given by the following. Here \(a\) is an arbitrary constant.\[\Large x^2-ay^2=r^2\]First off: What's the radius \(r\)? Second, translate it so that it's centered about your point. And third, use the given intersection point to find \(a\).

OpenStudy (anonymous):

The general form is \[ \frac{ (x-5)^2} {25}-\frac{(y+6)^2}{b^2}=1 \] Replace x by 15 and yby zerp and solve for b

OpenStudy (anonymous):

okay I see what your saying so far :)

OpenStudy (anonymous):

you get \[4-\frac{36}{b^2}=1\\ \frac{36}{b^2}=3\\ b^2 =12 \]

OpenStudy (anonymous):

did I have to graph any of this

OpenStudy (anonymous):

\[ \frac { (x-5)^2}{25}-\frac{(y+6)^2}{12} =1 \] No you do not need to graph it.

OpenStudy (anonymous):

okay the answer is (x-5)^2/25-(y+6)^2/12=1 and the center is hk which os -5 and 6